# Gorenstein singularities with $\mathbb{G}_{m}$-action and moduli spaces of holomorphic differentials

### v1

@@ Abstract.
Given a holomorphic differential on a smooth complex algebraic curve, we associate to it a Gorenstein curve singularity with $\mathbb{G}_{m}$-action via a test configuration. This construction decomposes the strata of holomorphic differentials with prescribed orders of zeros into negatively graded versal deformation spaces of such singularities, refining Pinkham’s correspondence between monomial curve singularities and Weierstrass semigroups to the case of Gorenstein singularities with multiple branches in the framework of Looijenga’s deformations with good $\mathbb{G}_{m}$-action. Additionally, this construction provides a natural description for the singular curves that appear in the boundary of the versal deformation spaces, generalizing various special cases from symmetric semigroups and local complete intersections to arbitrary Gorenstein curves that admit canonical divisors with prescribed orders of zeros.
Our construction provides a uniform approach to describe the resulting singularities and their invariants, such as weights and characters, initially studied by Alper–Fedorchuk–Smyth. As an application, we classify the unique Gorenstein singularity with $\mathbb{G}_{m}$-action for each nonvarying stratum of holomorphic differentials in the work of Chen–Möller and Yu–Zuo, identify each nonvarying stratum with the locus of smooth deformations of the corresponding singularity, and study when these nonvarying strata can be compactified by weighted projective spaces. Additionally, we classify such singularities with bounded $\alpha$-invariants in the Hassett–Keel log minimal model program for ¯$\overline{\mathcal{M}}_{g}$. We also study the slopes of these singularities and utilize them to bound the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$. Finally, we show that the loci of subcanonical points with fixed semigroups have trivial tautological rings and provide a criterion to determine whether they are affine.
Research of D.C. is supported by National Science Foundation Grant DMS-2301030, Simons Travel Support for Mathematicians, and a Simons Fellowship.
Research of F.Y. is supported by Fundamental Research Funds for the Central Universities 2024FZZX02-01-01.
@@ 1. Introduction
For $g,n\geq 1$, let $\mu=(m_{1},\ldots,m_{n})$ be a partition of $22g-$ 2, where $m_{i}\in\mathbb{N}$ and $\sum_{i=1}^{n}m_{i}=2g-2$. Denote by $\mathbb{P}\mathcal{H}(\mu)$ the moduli space parameterizing smooth and connected complex algebraic curves $C$ with distinct marked points $p_{1},\ldots,p_{n}$ such that $\sum_{i=1}^{n}m_{i}p_{i}$ is a canonical divisor. Equivalently, $\mathbb{P}\mathcal{H}(\mu)$ parameterizes holomorphic differentials $\omega\in H^{0}(C,K)$ (up to scaling) of zero type specified by $\mu$, where $K$ is the canonical bundle.
Since the union of $\mathbb{P}\mathcal{H}(\mu)$ for all partitions $\mu$ of $22g-$ 2 stratifies the (projectivized) Hodge bundle of holomorphic differentials, $\mathbb{P}\mathcal{H}(\mu)$ is called the (projectivized) stratum of holomorphic differentials of type $\mu$. For special $\mu$, $\mathbb{P}\mathcal{H}(\mu)$ can have up to three connected components, due to the spin parity and hyperelliptic structure; see [KZ03]. We use $\mathbb{P}\mathcal{H}(\mu)^{\ast}$ for $\ast\in\{\operatorname{odd},\operatorname{even},\operatorname{hyp},\operatorname{nonhyp}\}$ to denote these connected components of the strata. As a convention, a stratum in this paper will refer to a connected component if the corresponding stratum is disconnected. When $\mu$ contains many entries of the same value, we also use the exponential notation; e.g., $\mu=(1^{2g-2})$ is the signature of differentials with $22g-$ 2 simple zeros.
@@ 1.1. Gorenstein singularities with $\mathbb{G}_{m}$-action and test configurations
We remark that the special case of $n=1$ was studied in detail by Pinkham in [Pin74] (for unibranched singularities with $\mathbb{G}_{m}$-action that are not necessarily Gorenstein). In this case, $(Y,q)$ is a monomial singularity and $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ can be identified with the locus $\mathcal{M}_{g,1}^{H}\subset\mathcal{M}_{g,1}$ parameterizing smooth pointed curves $(C,p)$, where the Weierstrass semigroup $H$ of $p$ is generated by the exponents of the monomials defining $(Y,q)$. Therefore, our construction can be regarded as a refined generalization of Pinkham’s result from the case of unibranched singularities to Gorenstein singularities with multiple branches; see also [Loo84b, Appendix] for a formal description of such constructions with modular interpretations.
@@ 1.2. Uniqueness of singularities for the nonvarying strata
There is a unique isomorphism class of Gorenstein singularities with $\mathbb{G}_{m}$-action for every nonvarying stratum $\mathbb{P}\mathcal{H}(\mu)$ of holomorphic differentials of type $\mu$ in the following list: $\mu=(4)^{\operatorname{odd}}$, $(3,1)$, $(2,2)^{\operatorname{odd}}$, $(2,1,1)$, $(6)^{\operatorname{odd}}$, $(6)^{\operatorname{even}}$, $(5,1)$, $(4,2)^{\operatorname{even}}$, $(4,2)^{\operatorname{odd}}$, $(3,3)^{\operatorname{nonhyp}}$, $(3,2,1)$, $(2,2,2)^{\operatorname{odd}}$, $(6,2)^{\operatorname{odd}}$, $(5,3)$, as well as the hyperelliptic strata $(2g-2)^{\operatorname{hyp}}$ and $(g-1,g-1)^{\operatorname{hyp}}$, where each of these nonvarying strata can be identified with the locus of smooth deformations in the versal deformation space of the corresponding singularity.
@@
Indeed, we can describe the defining equations for all the above singularities as well as their weighted projective embeddings; see Section 5. The two possible exceptional cases in the second part of Theorem 1.2 are due to the fact that the corresponding singularities have nonzero infinitesimal obstructions; see Remark 5.3. For the remaining cases, with the defining equations of the singularities at hand, we can also describe explicitly the weighted projective spaces that compactify the corresponding strata; see Example 5.2 for a sample description.
@@ 1.3. Weights and characters of singularities
As mentioned in Theorem 1.1, the weights and characters of the $\mathbb{G}_{m}$-action on the space of differentials of $Y$ can be determined explicitly through our construction. These important invariants can help us better understand the Hassett–Keel log minimal model program for ¯$\overline{\mathcal{M}}_{g}$, as studied extensively by Alper–Fedorchuk–Smyth in [AFS16]. By utilizing our construction, we can systematically compute these invariants, paving an avenue towards a further classification of related singularities. To this end, we provide various examples and demonstrate the computations in Section 3. Moreover, in Section 4, we establish the following result.
@@ 1.4. Singularities with bounded $\alpha$-invariants
If a new singularity first appears in the log minimal model program for ¯$\overline{\mathcal{M}}_{g}$ with respect to the divisor class $K+\alpha\delta$, where $K$ is the canonical class and $\delta$ is the total boundary divisor class of ¯$\overline{\mathcal{M}}_{g}$, then $\alpha$ is called the $\alpha$-invariant of the singularity. There is a critical value $\alpha=3/8$ that plays a significant role in this process for large genus, as explained in [AFS16, Section 3.4]. Roughly speaking, the polarization on the GIT quotient of the Chow variety of canonical curves is asymptotically proportional to $K+\frac{3}{8}\delta$. Therefore, it is an interesting question to classify Gorenstein singularities with $\alpha$-invariants above $3/8$, as asked in [AFS16, Problem 3.15]. As another application of our method, we provide the following classification.
@@
    $\mu=(a,a,b,b)^{\operatorname{hyp}}$, $(2a,b,b)^{\operatorname{hyp}}$, $(2a,2b)^{\operatorname{hyp}}$, $(g-1,g-1)^{\operatorname{hyp}}$, $(2g-2)^{\operatorname{hyp}}$, $(2g-2,0)^{\operatorname{hyp}}$, and $(2g-2,0,0)^{\operatorname{hyp}}$, where the underlying curves are hyperelliptic, the zeros of even order are Weierstrass points, the zeros of equal order form hyperelliptic conjugacy pairs, and the ordinary marked points (zeros of order 00) can be any points.
@@
We are also able to describe the defining equations for all the singularities that appear in Theorem 1.4. Some of these singularities are well-studied in the literature, while it seems the others can be difficult to figure out without invoking our viewpoint. Just for the reader to get a feel, for example, the defining equations for the unique singularity corresponding to the nonvarying stratum $\mathbb{P}\mathcal{H}(6,2)^{\operatorname{odd}}$ are given by
@@
1.5. Slopes of singularities and effective divisors in ¯$\overline{\mathcal{M}}_{g}$
@@
where $\delta$ is the total boundary divisor class of ¯$\overline{\mathcal{M}}_{g}$ and $\lambda_{1}$ is the first Chern class of the Hodge bundle. In particular, if $B$ represents the class of a moving curve in ¯$\overline{\mathcal{M}}_{g}$, then the slope of $B$ can provide a lower bound for the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$ through their nonnegative intersection numbers. This is due to the fact that an effective divisor cannot contain the union of such moving curves.
Determining the lower bound for the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$ is an important, long-standing problem; see [CFM13] for a survey discussion. In our context, motivated by [AFS16, Theorem 4.2], we can similarly define the slope for singularity classes $(Y,q)$ that are associated to $\mathbb{P}\mathcal{H}(\mu)$. Moreover, if $\mathbb{P}\mathcal{H}(\mu)$ dominates $\mathcal{M}_{g}$ and the discriminant locus of singular deformations in ${\rm Def}^{-}(Y,q)$ consists of generically nodal curves, then we can use the slope of $(Y,q)$ to bound the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$.
@@
Let $\mu$ be a partition of $22g-$ 2 such that $\mu$ has at least $g-1$ positive entries and $\mu\neq(2^{g-1})^{\operatorname{even}}$. Let $(Y,q)$ be a singularity class associated to a generic canonical divisor of type $\mu$. If the discriminant locus of singular deformations in ${\rm Def}^{-}(Y,q)$ is a $\mathbb{Q}$-Cartier divisor parameterizing generically nodal curves, then the slope of $(Y,q)$ is a lower bound for the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$.

### v2

@@ Abstract.
Given a holomorphic differential on a smooth complex algebraic curve, we associate to it a Gorenstein curve singularity with $\mathbb{G}_{m}$-action via a test configuration. This construction decomposes the strata of holomorphic differentials with prescribed orders of zeros into negatively graded miniversal deformation spaces of such singularities. Additionally, it provides a natural description for the singular curves that appear in the boundary of the miniversal deformation spaces.
Our approach leads to a number of applications. We classify the unique Gorenstein singularity with $\mathbb{G}_{m}$-action for each nonvarying stratum of holomorphic differentials and study when these nonvarying strata can be compactified by weighted projective spaces. Moreover, extending the classical results about $ADE$ singularities, we establish the $K(\pi,1)$-property for non-hypersurface complete intersection singularities of type $U_{7}$, $U_{8}$, $U_{9}$, and $S_{k}$. We also study singularities with bounded $\alpha$-invariants in the log minimal model program for $\overline{\mathcal{M}}_{g}$ and utilize them to bound the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$. Finally, we show that the loci of subcanonical points with fixed semigroups have trivial tautological rings and provide a criterion to determine whether they are affine varieties.
2020 Mathematics Subject Classification:
Primary: 14H10; Secondary: 13H10, 14B07, 14H20, 14H45, 14H55, 32G15
Research of D.C. was supported by the National Science Foundation under Grant DMS-2301030, Simons Travel Support for Mathematicians, and a Simons Fellowship under Record ID SFI-MPS-SFM-00005694.
Research of F.Y. was supported by Fundamental Research Funds for the Central Universities 2024FZZX02-01-01.
@@ 1. Introduction
For $g,n\geq 1$, let $\mu=(m_{1},\ldots,m_{n})$ be a partition of 2g22g-2, where $m_{i}\in\mathbb{N}$ and $\sum_{i=1}^{n}m_{i}=2g-2$. Denote by $\mathbb{P}\mathcal{H}(\mu)$ the moduli space parameterizing smooth and connected complex algebraic curves $C$ with distinct marked points $p_{1},\ldots,p_{n}$ such that $\sum_{i=1}^{n}m_{i}p_{i}$ is a canonical divisor. Equivalently, $\mathbb{P}\mathcal{H}(\mu)$ parameterizes holomorphic differentials $\omega\in H^{0}(C,K)$ (up to scaling) of zero type specified by $\mu$, where $K$ is the canonical bundle.
Since the union of $\mathbb{P}\mathcal{H}(\mu)$ for all partitions $\mu$ of 2g22g-2 stratifies the (projectivized) Hodge bundle of holomorphic differentials, $\mathbb{P}\mathcal{H}(\mu)$ is called the (projectivized) stratum of holomorphic differentials of type $\mu$. For special $\mu$, the stratum $\mathbb{P}\mathcal{H}(\mu)$ can have up to three connected components, due to the spin parity and hyperelliptic structure; see [KZ03]. We use $\mathbb{P}\mathcal{H}(\mu)^{\ast}$ for $\ast\in\{\operatorname{odd},\operatorname{even},\operatorname{hyp},\operatorname{nonhyp}\}$ to denote these connected components of the strata. As a convention, a “stratum” in this paper will refer to a connected component if the corresponding $\mathbb{P}\mathcal{H}(\mu)$ is disconnected. When $\mu$ contains many entries of the same value, we also use the exponential notation; e.g., $\mu=(1^{2g-2})$ is the signature of differentials with 2g22g-2 simple zeros.
@@ 1.1. Gorenstein singularities with $\mathbb{G}_{m}$-action and test configurations
We remark that the special case of $n=1$ was studied in detail by Pinkham in [Pin74] (for unibranched singularities with $\mathbb{G}_{m}$-action which are not necessarily Gorenstein). In this case, $(Y,q)$ is a monomial singularity and $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ can be identified with the locus $\mathcal{M}_{g,1}^{H}\subset\mathcal{M}_{g,1}$ parameterizing smooth pointed curves $(C,p)$, where the Weierstrass semigroup $H$ of $p$ is generated by the exponents of the monomials defining $(Y,q)$. Therefore, our construction can be regarded as a refined generalization of Pinkham’s result from the case of unibranched singularities to Gorenstein singularities with multiple branches; see also [Loo84b, Appendix] for a formal description of such constructions with modular interpretations.
@@ 1.2. Uniqueness of singularities for the nonvarying strata
There is a unique isomorphism class of Gorenstein singularities with $\mathbb{G}_{m}$-action for every nonvarying stratum $\mathbb{P}\mathcal{H}(\mu)$ of holomorphic differentials of type $\mu$ in the following list: $\mu=(4)^{\operatorname{odd}}$, $(3,1)$, $(2,2)^{\operatorname{odd}}$, $(2,1,1)$, $(6)^{\operatorname{odd}}$, $(6)^{\operatorname{even}}$, $(5,1)$, $(4,2)^{\operatorname{even}}$, $(4,2)^{\operatorname{odd}}$, $(3,3)^{\operatorname{nonhyp}}$, $(3,2,1)$, $(2,2,2)^{\operatorname{odd}}$, $(6,2)^{\operatorname{odd}}$, $(5,3)$, as well as the hyperelliptic strata $(2g-2)^{\operatorname{hyp}}$ and $(g-1,g-1)^{\operatorname{hyp}}$, where each of these nonvarying strata can be identified with the locus of smooth deformations in the miniversal deformation space of the corresponding singularity.
@@
Finally, there is a unique isomorphism class of Gorenstein singularities with $\mathbb{G}_{m}$-action associated to every locus of hyperelliptic differentials with signature
$\mu=(2a_{1},\ldots,2a_{m},b_{1},b_{1},\ldots,b_{n},b_{n}),$
where each zero of order $2a_{i}$ is a Weierstrass point and each pair of zeros of order $b_{j}$ is hyperelliptic conjugate.
Indeed, for the above nonvarying strata in low genus, we can describe the defining equations for the corresponding singularities as well as their weighted projective embeddings; see Section 5. The two possible exceptional cases in the second part of Theorem 1.2 are due to the fact that the corresponding singularities have nonzero infinitesimal obstructions; see Remark 5.3. For the remaining cases, with the defining equations of the singularities at hand, we can also describe explicitly the weighted projective spaces that compactify the corresponding strata; see Example 5.2 for a sample description.
1.3. Adding ordinary marked points and the $K(\pi,1)$-property
Recall that a space is $K(\pi,1)$ if its (orbifold) universal cover is contractible; equivalently, if its homotopy groups of degree higher than one all vanish. The geometry of $\mathbb{P}\mathcal{H}(\mu)$ and its decomposition into deformation spaces $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ can motivate a number of intriguing $K(\pi,1)$ and affine patching related questions; see [Kon97, LM14, Mon17, Gia24, Gia25, Qiu25] for some related conjectures and results.
Note that our result has identified each of the nonvarying strata with the locus of smooth deformations of the corresponding singularity, which can be regarded as the discriminant hypersurface complement in the miniversal deformation space. A conjecture (sometimes attributed to Saito; see [Arn93, p. 26] and [Loo84a, p. 185]) says that the discriminant complement in the miniversal deformation space of a positive-dimensional isolated complete intersection singularity is $K(\pi,1)$. The cases for $A$ and $DE$ singularities were established by Brieskorn and Deligne, respectively; see [Bri73, Del72] as well as [Loo84a, Chapter 9] for a more detailed discussion.
Although the original definition of nonvarying strata was stated for $\mu$ with only positive entries, we can also add entries of 0 into $\mu$ which amounts to adding ordinary marked points away from the original zeros. From the perspective of associated singularities, the former and latter determine each other; see Section 5.14 for a detailed description. This observation has a striking application for the $K(\pi,1)$-property of certain non-hypersurface complete intersection singularities as follows.
Theorem 1.3.
Let $(Y^{\prime},q^{\prime})$ be a singularity of type $(\mu,0)$ obtained by adding an ordinary marked point away from the zeros in the construction of a singularity $(Y,q)$ of type $\mu$. If $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ is $K(\pi,1)$, then $\mathbb{P}{\rm Def}^{-}_{s}(Y^{\prime},q^{\prime})$ is also $K(\pi,1)$.
In particular, the $K(\pi,1)$-property holds for the discriminant complement in the miniversal deformation space of the singularities $U_{7}$, $U_{8}$, $U_{9}$, $S_{2g+2}$, and $S_{2g+3}$, which arise from the nonvarying strata for $\mu^{\prime}=(4,0)^{\operatorname{odd}}$, $(3,1,0)$, $(6,0)^{\operatorname{even}}$, $(2g-2,0,0)^{\operatorname{hyp}}$, and $(g-1,g-1,0,0)^{\operatorname{hyp}}$, respectively.
We refer to Theorem 5.4 for a detailed description of the above singularities. See also Section 5.15 for a related discussion about the $K(\pi,1)$-property of singularities associated to hyperelliptic differentials. Additionally, some of the nonvarying strata in our classification correspond to simple space curve singularities of type $T_{7}$ (for $\mu=(2,2)^{\operatorname{odd}}$), $T_{8}$ (for $\mu=(2,1,1)$), $T_{9}$ (for $\mu=(4,2)^{\operatorname{even}}$), $W_{8}$ (for $\mu=(6)^{\operatorname{odd}}$), $W_{9}$ (for $\mu=(5,1)$), $Z_{9}$ (for $\mu=(3,3)^{\operatorname{nonhyp}}$), and $Z_{10}$ (for $\mu=(8)^{\operatorname{even}}$); see also [Loo84a, (7.22)] and [Ste15, Table 2]. Therefore, the $K(\pi,1)$-conjecture for them is thus equivalent to that of the corresponding strata of differentials.
@@ 1.4. Weights and characters of singularities
As mentioned in Theorem 1.1, the weights and characters of the $\mathbb{G}_{m}$-action on the space of differentials of $Y$ can be determined explicitly through our construction. These important invariants can help us better understand the Hassett–Keel log minimal model program for $\overline{\mathcal{M}}_{g}$, as studied extensively by Alper–Fedorchuk–Smyth in [AFS16]. By utilizing our construction, we can systematically compute these invariants, paving an avenue towards a further classification of related singularities. To this end, we provide various examples and demonstrate the computations in Section 3. Moreover, in Section 4, we establish the following result.
@@ 1.5. Singularities with bounded $\alpha$-invariants
If a new singularity first appears in the log minimal model program for $\overline{\mathcal{M}}_{g}$ with respect to the divisor class $K+\alpha\delta$, where $K$ is the canonical class and $\delta$ is the total boundary divisor class of $\overline{\mathcal{M}}_{g}$, then $\alpha$ is called the $\alpha$-invariant of the singularity. There is a critical value $\alpha=3/8$ that plays a significant role in this process for large genus, as explained in [AFS16, Section 3.4]. Roughly speaking, the polarization on the GIT quotient of the Chow variety of canonical curves is asymptotically proportional to $K+\frac{3}{8}\delta$. Therefore, it is an interesting question to classify Gorenstein singularities with $\alpha$-invariants above $3/8$, as asked in [AFS16, Problem 3.15]. As another application of our method, we provide the following classification.
@@
    $\mu=(a,a,b,b)^{\operatorname{hyp}}$, $(2a,b,b)^{\operatorname{hyp}}$, $(2a,2b)^{\operatorname{hyp}}$, $(g-1,g-1)^{\operatorname{hyp}}$, $(2g-2)^{\operatorname{hyp}}$, $(2g-2,0)^{\operatorname{hyp}}$, and $(2g-2,0,0)^{\operatorname{hyp}}$, where the underlying curves are hyperelliptic, the zeros of even order are Weierstrass points, the zeros of equal order form hyperelliptic conjugacy pairs, and the ordinary marked points (zeros of order 0) can be any points.
@@
We are also able to describe the defining equations for all the singularities that appear in Theorem 1.5. Some of these singularities are well-studied in the literature, while it seems the others can be difficult to figure out without invoking our perspective. Just for the reader to get a feel, for example, the defining equations for the unique singularity corresponding to the nonvarying stratum $\mathbb{P}\mathcal{H}(6,2)^{\operatorname{odd}}$ are given by
@@
1.6. Slopes of singularities and effective divisors in $\overline{\mathcal{M}}_{g}$
@@
where $\delta$ is the total boundary divisor class of $\overline{\mathcal{M}}_{g}$ and $\lambda_{1}$ is the first Chern class of the Hodge bundle. In particular, if $B$ represents the class of a moving curve in $\overline{\mathcal{M}}_{g}$, then the slope of $B$ can provide a lower bound for the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$ through their intersection numbers. This is due to the fact that an effective divisor $D$ cannot contain the union of such moving curves, and hence $D$ and $B$ must have a nonnegative intersection number.
Determining the lower bound for the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$ is an important, long-standing problem; see [CFM13] for a survey discussion. In our context, motivated by [AFS16, Theorem 4.2], we can similarly define the slope for singularity classes $(Y,q)$ that are associated to $\mathbb{P}\mathcal{H}(\mu)$. Moreover, if $\mathbb{P}\mathcal{H}(\mu)$ dominates $\mathcal{M}_{g}$ and the discriminant locus of singular deformations in ${\rm Def}^{-}(Y,q)$ consists of generically nodal curves, then we can use the slope of $(Y,q)$ to bound the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$.
@@
Let $\mu$ be a partition of 2g22g-2 such that $\mu$ has at least g1g-1 positive entries and $\mu\neq(2^{g-1})^{\operatorname{even}}$. Let $(Y,q)$ be a singularity class associated to a generic canonical divisor of type $\mu$. If the discriminant locus of singular deformations in ${\rm Def}^{-}(Y,q)$ is a $\mathbb{Q}$-Cartier divisor parameterizing generically nodal curves, then the slope of $(Y,q)$ is a lower bound for the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$.